OpenStudy (anonymous):

Multiply. (-2t)^2 (4t^3)^2

OpenStudy (bahrom7893):

4t^2 * 16t^6

OpenStudy (bahrom7893):

= 64t^8

OpenStudy (bahrom7893):

Please click on become a fan if I helped, I really want to get to the next level!! Thanks =)

OpenStudy (anonymous):

hahaha, i've been forgetting to ask people to do that. maybe thats why ive stopped getting fans 0.0

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

help me with my next one mathfreak and I'll fan ya

OpenStudy (anonymous):

ok, go right ahead :)

OpenStudy (anonymous):

^18sqrt(-6)^18

OpenStudy (anonymous):

can you rewrite that. why is there a "^" infront of the equation?

OpenStudy (bahrom7893):

what?

OpenStudy (anonymous):

Not sure how to describe it but it isn't the big number in front of the square root but a little number like a power...make sense?

OpenStudy (bahrom7893):

okay i'll let mathfreak deal with this one..

OpenStudy (bahrom7893):

can you take a picture?

OpenStudy (bahrom7893):

oh I know

OpenStudy (bahrom7893):

it means 18th root of

OpenStudy (bahrom7893):

see it means 18th root of (-6)^18

OpenStudy (bahrom7893):

or: (-6)^[(18)*(1/18)]

OpenStudy (anonymous):

yes that is it...LOL

OpenStudy (bahrom7893):

-6^(18/18) = -6^1 = -6

OpenStudy (bahrom7893):

you gotta fan me again.. lol pity u can't

OpenStudy (anonymous):

LOL....."Clapping fro bahrom7893" since I can fan again!

OpenStudy (bahrom7893):

LOL can't and for, i believe..

OpenStudy (anonymous):

or spell??? is there spelling lessons in here too....double bonus

OpenStudy (anonymous):

sorry, i left to see my own question. I was being selfish. hey, who said that was bad? i have Ayn Rand on my side XP

OpenStudy (anonymous):

okay so solve this mathfreak...5x^2=35 and find the x intercepts

OpenStudy (anonymous):

+or - square root of 7

OpenStudy (anonymous):

I think the answer is +/-sqrt7?

OpenStudy (anonymous):

5x^2=35 x^2=35/7 x^2= 5 x=+/-sqrt7

OpenStudy (anonymous):

YES! I got it right then! Thank you, you are now my fan!

OpenStudy (anonymous):

yesh! thanks!